# GaINeR: Geometry-Aware Implicit Neural Representation for Image Editing

### v1

@@ 5 Conclusion
Our framework bridges coordinate-based MLPs with explicit geometric priors, yielding a unified model capable of high-fidelity reconstruction and photometrically consistent transformations. Comprehensive experiments demonstrate that GaINeR surpasses prior implicit representations such as SIREN, WIRE, and MiraGe in both quantitative accuracy and visual quality, while offering greater interpretability through its Gaussian-based encoding.
@@
6 Acknowledgment
The work of W. Jakubowska and P. Spurek was supported by the National Centre of Science (Poland) Grant No. 2023/50/E/ST6/00068. M. Zieliński and D. Belter were supported by the National Science Centre, Poland, under research project no UMO-2023/51/B/ST6/01646. The scientific work was carried out using the infrastructure of the Poznan Supercomputing and Networking Center.
Supplementary materials:
GaINeR: Geometry-Aware Implicit Neural Representation for Image Editing
Weronika Jakubowska*1,† Mikołaj Zieliński*2 Rafał Tobiasz*3,4 Krzysztof Byrski4
Maciej Zięba1 Dominik Belter2 Przemysław Spurek3,4
1Wrocław University of Science and Technology  2Poznań University of Technology
3IDEAS Research Institute 4Jagiellonian University
weronika.jakubowska@pwr.edu.pl
*Equal contribution
7 Extended Experimental Analysis
This section presents additional experiments that complement the main results. It begins by outlining the difficulties encountered by prompt-based LLM and diffusion editors when performing controlled geometric modifications. This is followed by further editing examples, physics-inspired manipulations, and an ablation study. Together, these extended results provide a broader view of the method’s capabilities and behavior.
7.1 Limitations of Prompt-Based LLM and Diffusion Editors
Recent LLM- and diffusion-based editors enable high-level semantic modifications, but they lack geometric controllability and often introduce unpredictable artifacts when altering object shape, scale, or structure. Because these models operate through generative resampling, they do not preserve underlying geometry and offer limited support for continuous, user-defined transformations or temporally stable animation. As illustrated in Fig. 10, methods such as GPT-IMAGE-1 [gtp_image_1], Flux [labs2025flux1kontextflowmatching] and Seedream 4.0 [seedream2025seedream40nextgenerationmultimodal] fail to perform even simple geometric edits, consistently producing shape distortions or hallucinated content. In contrast, our INR-based representation provides a continuous, function-level encoding that enables precise and deterministic geometric edits including bending, stretching, and scaling without altering object identity or introducing hallucinated content.
Figure 10: Comparison of geometric edit performance across models. Starting from a generated image of a cat, each method was asked to apply the prompt: “Enlarge only the cat’s left ear by about 150%, keeping the rest of the cat completely unchanged.” For our method, the edit was applied directly through our representation using a controlled geometric transformation implemented in Blender. GaINeR method produces a clean, targeted edit while keeping the cat’s appearance intact. In contrast, Flux cuts off part of the ear, Seedream enlarges the ear but also scales the whole cat, and GPT-IMAGE-1 makes the ear much larger than requested. This shows that prompt-based generative editors struggle with simple geometric edits.
7.2 Error Comparisons
@@
7.3 Editing Comparisons
In Fig. 15, we provide additional qualitative comparisons illustrating how the models respond to a broader range of manual geometric modifications. These examples primarily involve bending and, in some cases, moderate stretching performed in Blender [blender]. Across most scenarios, the GaINeR maintains stable geometry and appearance, producing visually coherent results even under stronger or more irregular deformations. MiraGe [waczynskamirage] occasionally exhibits structural inconsistencies under such modifications; however, in several examples both approaches perform comparably well.
To further illustrate the flexibility of the editing process, Fig. 12 shows an example in which the extracted object is manually modified and subsequently placed into a newly generated background. This workflow enables precise, user-controlled adjustments while preserving consistency between the modified object and the environment.
To further assess the model’s behavior under simple geometric edits, we change the scale of the main object in the image by either enlarging or shrinking it. As shown in Fig. 13, these operations do not introduce visual artifacts, and the object remains structurally continuous and visually consistent in both cases.
Beyond static manual edits, the editing pipeline also supports generating animations in Blender, enabling controlled simulations of various object motions. By applying frame-wise geometric transformations, we can assess the temporal stability of the reconstruction process. As illustrated in Fig. 14, GaINeR maintains structural coherence and produces artifact-free frames throughout the animated sequence, demonstrating its robustness under continuous, motion-induced transformations.
Figure 12: Object extraction, manual editing in Blender, and background replacement produced using GaINeR. The edited object is reintegrated into a newly generated scene while maintaining visual consistency.
Figure 13: GaINeR tested under scale-based geometric modifications, including object enlargement and reduction. In all cases, the method yields stable and visually coherent results.
@@ 7.3 Editing Comparisons
Figure 15: Extended qualitative assessment of MiraGe [waczynskamirage] and GaINeR under manually introduced geometric transformations. Both models handle the fisherman, puffin, and statue scenes equally well. In contrast, MiraGe shows artifacts in several cases: around the woman’s abdomen, near the elephant’s neck when using the mirror camera, and along the body of the lizard and the neck of the gray crowned crane under the same camera setting.
7.4 Physics Evaluations
We further investigate the dynamic capabilities of GaINeR through three distinct simulation scenarios powered by the Taichi Elements engine [taichi_elements]. These experiments demonstrate the versatility of our explicit Gaussian embeddings when utilized as physical particles.
@@ 7.4 Physics Evaluations
Finally, we explore complex multi-material interactions within a single scene (Fig. 16, bottom). Leveraging the explicit spatial coordinates of GaINeR, we segment a single trained image representation into distinct regions based on pixel position. Gaussians corresponding to the text were assigned granular (sand) material properties, while the upper region was assigned fluid (water) properties. This confirms that GaINeR allows for spatially-aware semantic assignment of physical parameters, enabling diverse materials to interact seamlessly within a unified simulation.
@@
Figure 16: Visualizations of physics-based simulations. Top: An elastic apple deforming under constant gravity upon impact. Middle: A bouquet subjected to oscillating lateral forces, demonstrating structural flexibility. Bottom: A multi-material simulation showcasing the interaction between a granular object and a fluid volume, demonstrating complex mixing and flow dynamics.
7.5 Object Integration and Compositing
We compare the workflow for integrating edited foreground objects with backgrounds in GaINeR versus MiRaGe [waczynskamirage].
Based on the collage capabilities presented in [waczynskamirage], MiRaGe typically handles multi-object integration by arranging Gaussian primitives within a shared 3D space. Consequently, the final image is produced via a single rasterization pass. This approach relies heavily on the splatting engine to handle depth sorting and occlusion. However, as observed in the qualitative comparisons in our main paper (see physics simulation figures), this approach can lead to unintended transparency artifacts. Since the object is represented as a cloud of discrete primitives without a continuous surface mesh, deformation (such as stretching) can create gaps between Gaussians or reduce accumulated density, causing the background to bleed through the foreground object.
In contrast, our framework adopts a layer-based compositing workflow. We learn a precise, continuous alpha mask of the object alongside its RGB appearance. The modified object is rendered together with its mask to produce explicit color and alpha channels. The final scene is assembled by alpha-blending the rendered foreground over the inpainted background.
By explicitly separating foreground reconstruction from background integration, GaINeR ensures crisp object boundaries and maintains solid opacity. Even under significant deformation, the learned alpha mask ensures the object fully occludes the background, eliminating the "ghosting" or transparency artifacts observed in the baseline methods.
7.6 Size and Time Comparisons
We compare the computational characteristics of GaINeR with several representative baselines, reporting training time, inference throughput, and model size. To ensure consistency across methods with differing input-size constraints, all models were trained on the DIV2K dataset [Agustsson_2017_CVPR_Workshops] subjected to $3\times$ bicubic downscaling. Training was performed on an NVIDIA RTX 5090 GPU, except for MiraGe [waczynskamirage], which requires older CUDA support and was therefore trained on an NVIDIA RTX 3090. For GaINeR, we report only the components required at inference (excluding optimizers, training buffers, and hash-grid encoders) so that the model size reflects the minimal representation used for rendering.
The results in Tab. 3 expose a clear trade-off between editability, reconstruction quality, and runtime/storage costs. Only MiraGe and GaINeR provide native editability. These two methods also achieve the highest PSNR in our experiments, indicating that editability in our setting did not come at the expense of reconstruction fidelity. However, this capability incurs a computational cost: both editable models have substantially larger serialized footprints and slower end-to-end performance compared to lightweight baselines (SIREN [sitzmann2020implicit], FINER [liu2024finer], Gauss [ramasinghe2022beyond]) and extremely compact pipelines (GaussianImage [zhang2024gaussianimage]). Notably, our primary editable baseline, MiraGe, is more than twice as large as our model and, due to its densification process, does not maintain a constant representation size during training, whereas our method preserves a fixed and compact footprint throughout.
Model Time Size [MB]
Train [s] Inference [FPS]
WIRE [saragadam2023wire] 410 34.01 0.76
SIREN [sitzmann2020implicit] 89 163.90 0.76
FINER [liu2024finer] 122 83.84 0.76
Gauss [ramasinghe2022beyond] 134 105.88 0.76
GaussianImage [zhang2024gaussianimage] 236 5187.82 0.074
MiraGe [waczynskamirage] 299 358.07 109.82
GaINeR (our) 792 2.50 51.43
7.7 Ablation study

### v2

@@ 5 Conclusion
Our framework bridges coordinate-based MLPs with explicit geometric priors, yielding a unified model capable of high-fidelity reconstruction and photometrically consistent transformations. Comprehensive experiments demonstrate that GaINeR surpasses prior implicit representations such as SIREN, WIRE, and MiraGe in both quantitative accuracy and visual quality, while offering greater interpretability through its Gaussian-based encoding. In addition, the continuous formulation of GaINeR naturally supports arbitrary-resolution image synthesis, enabling seamless super-resolution and upscaling without retraining.
@@
Limitations While GaINeR enables 3D lifting, its geometry is bounded by the quality of the off-the-shelf monocular depth estimator. Errors in depth prediction can lead to geometric artifacts.
Supplementary Material
Weronika JakubowskaMikołaj ZielińskiRafał TobiaszKrzysztof ByrskiMaciej ZiębaDominik BelterPrzemysław Spurek
Appendix 0.A Extended Experimental Analysis
This section presents additional experiments and analyses that complement the main results. Specifically, we provide detailed error comparisons, further editing examples, physics-inspired evaluations, novel 2D-to-3D transformations, computational efficiency comparisons (size and time) and an extensive ablation study. Together, these extended results provide a broader view of the method’s capabilities and behavior.
0.A.1 Error Comparisons
@@
0.A.2 Editing Comparisons
In Fig. 12, we provide additional qualitative comparisons illustrating how the models respond to a broader range of manual geometric modifications. Across most scenarios, GaINeR maintains stable geometry and appearance, producing visually coherent results even under stronger or more irregular deformations. MiraGe [waczynskamirage] exhibits structural inconsistencies under such modifications.
Figure 12: Extended qualitative assessment of MiraGe [waczynskamirage] and GaINeR under manually introduced geometric transformations. MiraGe shows artifacts in several cases: around the woman’s abdomen, near the elephant’s neck when using the mirror camera, and along the body of the lizard and the neck of the gray crowned crane under the same camera setting.
Beyond static manual edits, the editing pipeline also supports generating animations, enabling controlled simulations of various object motions. By applying frame-wise geometric transformations, we can assess the temporal stability of the reconstruction process. As illustrated in Fig. 13, GaINeR maintains structural coherence and produces artifact-free frames throughout the animated sequence, demonstrating its robustness under continuous, motion-induced transformations.
@@ 0.A.2 Editing Comparisons
0.A.3 Physics Evaluations
We further investigate the dynamic capabilities of GaINeR through two distinct simulation scenarios powered by the Taichi Elements engine [taichi_elements]. These experiments demonstrate the versatility of our explicit Gaussian embeddings when utilized as physical particles.
@@ 0.A.3 Physics Evaluations
Figure 14: Visualizations of physics-based simulations. Top: An elastic apple deforming under constant gravity upon impact. Middle: A bouquet subjected to oscillating lateral forces, demonstrating structural flexibility.
0.A.4 2D-to-3D Transformations
In the main manuscript, we established the capability of our framework to lift implicit 2D image representations into navigable 3D environments. Here, we provide additional qualitative results to further validate the robustness of this approach.
Fig. 15 showcases the 2D-to-3D lifting process applied to complex, isolated objects. The novel view renders demonstrate that GaINeR successfully extrudes flat 2D topology into coherent 3D volumes. Importantly, it preserves intricate geometric details such as overlapping petals and complex contours while maintaining strict photometric consistency without introducing severe stretching artifacts.
Furthermore, Fig. 16 illustrates the versatility of our framework on full-frame scenes, including classical artworks and dense landscapes. As the camera trajectory deviates from the original frontal viewpoint, the synthesized renders exhibit a highly realistic parallax effect. These results confirm that our model maintains textural coherence and correct spatial relationships across varying depth planes, all derived entirely from a single unposed image.
Figure 15: Qualitative results of 2D-to-3D lifting on isolated objects. By leveraging monocular depth maps to spatially translate the learned Gaussian embeddings along the Z-axis, GaINeR accurately extrudes flat 2D representations into coherent 3D volumes.
Figure 16: Immersive 2D-to-3D lifting on full-frame paintings and landscapes. Driven by the corresponding continuous depth gradients, GaINeR transforms traditional flat artworks into navigable volumetric scenes.
0.A.5 Ablation study
@@
0.A.6 Size and Time Comparisons
We compare the computational characteristics of GaINeR with several representative baselines, reporting training time, inference throughput, and model size. To ensure consistency across methods with differing input-size constraints, all models were trained on the DIV2K dataset [Agustsson_2017_CVPR_Workshops] subjected to $3\times$ bicubic downscaling. For GaINeR, we report only the components required at inference (excluding optimizers, training buffers, and hash-grid encoders) so that the model size reflects the minimal representation used for rendering.
The results in Table 4 expose a clear trade-off between editability, reconstruction quality, and runtime/storage costs. Only MiraGe and GaINeR provide native editability. These two methods also achieve the highest PSNR in our experiments, indicating that editability in our setting did not come at the expense of reconstruction fidelity. However, this capability incurs a computational cost: both editable models have substantially larger serialized footprints and slower end-to-end performance compared to lightweight baselines (SIREN [sitzmann2020implicit], FINER [liu2024finer], Gauss [ramasinghe2022beyond]) and extremely compact pipelines (GaussianImage [zhang2024gaussianimage]). Notably, our primary editable baseline, MiraGe, is more than twice as large as our model and, due to its densification process, does not maintain a constant representation size during training, whereas our method preserves a fixed and compact footprint throughout.
Model Time Size [MB]
Train [s] Inference [FPS]
WIRE [saragadam2023wire] 410 34.01 0.76
SIREN [sitzmann2020implicit] 89 163.90 0.76
FINER [liu2024finer] 122 83.84 0.76
Gauss [ramasinghe2022beyond] 134 105.88 0.76
GaussianImage [zhang2024gaussianimage] 236 5187.82 0.074
MiraGe [waczynskamirage] 299 358.07 109.82
GaINeR (our) 792 2.50 51.43